A bag has 4 red and 6 blue balls. One ball is picked at random. P( not red) =
- 32
- 35
- 25
- 710
Answer
B) 35
P( not red) = 1 - P( red) = 1 - 25 = 35. SLO M-09-C-04 · Knowledge
Mathematics MCQ · More Probability MCQs · Practise
B) 35
P( not red) = 1 - P( red) = 1 - 25 = 35. SLO M-09-C-04 · Knowledge
Mathematics MCQ · More Probability MCQs · Practise